### v1

# Towards Terminology Management Automation for Arabic

Abstract
This paper presents a method and supporting tools for automation of terminology management for Arabic. The tools extract lists of parallel terminology matching terms in foreign languages to their Arabic counterparts from field specific texts. This has significant implications as it can be used to improve consistent translation and use of terms in specialized Arabic academic books, and provides automated aid for enhancing cross-lingual text processing. This automation of terminology management aims to reduce processing time, and ensure use of consistent and correct terminology.
The extraction takes advantage of naturally occurring term translations. It considers several candidate phrases of varying lengths that co-occur next to the foreign terms.
Then it computes several similarity metrics, including lexicographic, phonetic, morphological, and semantic ones to decide the problem. We experiment with heuristic, machine learning, and ML with post processing approaches. This paper reports on a novel curated dataset for the task, an existing expert reviewed industry parallel corpora, and on the performance of the three approaches. The best approach achieved $94.9\%$ precision and $92.4\%$ recall.
Towards Terminology Management Automation for Arabic
Mahdi Nasser  and Laura Sayyah  and Fadi A. Zaraket Arab Center for Research and Policy Studies, Doha {mnasser,lsayyah,fzaraket} @dohainstitute.edu.qa
1 Introduction
Translation of academic literature is essential to knowledge transfer. Linguistic and structural complexities exist. These affect consistency and clarity of translated material and may hinder knowledge accessibility. Terminology management concerns the complexities involved in (i) specifying the term that should be used upon the introduction of a new concept to a language, and (ii) and enforcing the use, hence (iii) ensuring consistency and correctness.
The first and most important step is to build TermBase, a large database of parallel terms. Authors can utilize TermBase to enhance their writing processes when faced with a new concept; be it a foreign term to be translated to Arabic.
Recent advances in natural language processing and computational linguistics have paved the way for automated systems that assist in cross-lingual text processing. However, existing systems, including LLMs, often struggle with the nuances of terminology in cross-lingual specific expert fields, where precision is paramount. This paper introduces a novel methodology and a supporting framework to extract terminology mappings from existing Arabic books, a rich source of terms when building a TermBase.
When a foreign term occurs in Arabic books, usually between parentheses, it is preceded by its Arabic translation. Two problems arise. The exact scope of the Arabic term is ambiguous, and often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones. Editors and reviewers, who are usually expert researchers, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
Our work helps automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
The problem is formulated as follows: given a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$, denote candidate terms as the set of strings $s_{i}=w_{1}\ldots w_{i}$ ending before term $f$, $1\leq i\leq n$ as a candidate term. The goal is to identify the target candidate term $s_{t}$ best matching $f$. Our tool extracts linguistic features for each candidate match $(s_{i},f)$. The tool interpolates the features and aggregates their values into a resulting score that ranks the candidate terms.
By addressing the challenges of terminology mapping, this research moves us closer to a more integrated community where language barriers are significantly reduced, facilitating a greater exchange of knowledge and ideas across languages and cultures.
We make verified parts of the constructed dataset available online for the research community. To the best of our knowledge, this approach to TermBase extraction for Arabic is novel, and no comparable datasets currently exist.
1.1 Motivating example
Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme with different translations.
  • Context 1:
    - (l’ethnocentrisme) , : ¡(…) نّ ¿.
    Gloss: "His sharp criticism of ethnocentrism (l’ethnocentrisme) does not turn into cultural relativism, but rather into rational universality."
  • Context 2:
    اً (l’ethnocentrisme) ¡ ¿, ذّ ¡ ¿.
    Gloss: "He strongly criticized the ethnocentric tendency (l’ethnocentrisme) that seeks to replace primitive human ideas about the world and society with our European ideas."
Candidates for Context 1:
  • - Gloss: "And his sharp criticism of ethnocentrism does not turn into…"
  • - Gloss: "His sharp criticism of ethnocentrism turns into…"
  • - Gloss: "His sharp criticism of ethnocentrism"
  • - Gloss: "The sharp ethnocentrism"
  • -    (target term) Gloss: "Ethnocentrism"
  • Gloss: "Ethnicity"
Candidates for Context 2:
  • اً Gloss: "He strongly criticized the ethnocentric tendency"
  • اً Gloss: "The strong ethnocentric tendency"
  • Gloss: "The ethnocentric tendency"
  • (target term) Gloss: "Ethnocentrism"
  • Gloss: "Centralism"
The two contexts and the extracted translation candidate terms showcase the inconsistency across contexts, and also expose potential adoption of different candidates by readers of each context. Thus amplifying the problem.
This paper makes the following contributions:
  • We propose a novel method for terminology mapping that leverages naturally occurring term translations, combining NLP and data driven techniques.
  • We create a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
  • We evaluate and compare three approaches: heuristic, machine learning, and machine learning with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges. The results are impressive and convincing for deployment. The final model runs locally, requires no LLM interaction, and thus requires taking no IP risks from Arabic publishing houses.
We discuss TermBase creation in Section 2, the methodology in Section 3, briefly cover related work in  4, present and discuss our results in Section 5. We then discuss future directions and conclude.
2 Dataset creation
The source dataset is derived specifically from books published by our publishing house. These books span multiple academic fields, offering a diverse range of terminologies. The dataset includes terms in foreign languages (usually enclosed in parentheses) alongside their surrounding Arabic text.
The labeling process involved extracting candidate Arabic terms preceding each foreign term and annotating them with a binary label: True if the candidate is the correct translation of the foreign term, and False otherwise. To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed to correct the errors.
2.1 Source data selection
To ensure diversity and relevance, we selected texts from multiple domains. The candidate terms were extracted by identifying all strings of words, with a certain relative length with respect to the foreign term, immediately preceding a parenthetical foreign term. This ensures that potential translations are almost always captured for further analysis.
2.2 Annotated dataset construction
Each extracted candidate term, along with its associated foreign term, was presented as an individual instance. The features of this instance, extracted using NLP tools, are described in section 3.
Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 5.1 to facilitate the labeling process. This draft was then reviewed to ensure the accuracy of the labels.
3 Methodology
[…]
Figure 1: An overview of the process
Figure 1 illustrates the flow of our methodology. First, the tool extracts the candidates from a context (usually a paragraph) with a foreign term. Then, the feature vector of each candidate is computed. Finally, each vector is fed to a model which outputs a score that helps in choosing the target translation of the source foreign term.
3.1 Feature extraction
Given a candidate $s_{i}$, its foreign term $f$, and its context $c$, we extract features of the candidate that are going to be used later to predict which candidate is the target term. These features with their possible values are listed in Table 1 and explained in the following subsections.
Table 1: Features extracted for each candidate
Feature Value
Semantic similarity [0, 1]
Translation lexical
similarity [0, 1]
Transliteration lexical
similarity [0, 1]
Entity {PER, LOC, ORG, MISC}
POS {adj, adv, conj, misc, noun, noun_prop, part, prep, pron, verb}
Phonetic similarity {False, True}
3.1.1 Semantic similarity
We utilized the sentence transformer LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
After that, we compute semantic similarity as the cosine similarity between these 2 vectors Salton et al. (1975). This might seem sufficient but it’s not, which is due to the high similarity between candidates (given any two candidates, one of them is a substring of the other). Therefore, we need more features to handle peculiar cases.
Table 2: Semantic similarity example
Candidate Semantic similarity
Gloss: "Anne Herzberg" 0.635
Gloss: "Herzberg" 0.66
Table 2 shows how the semantic similarity alone score favors the $2^{nd}$ candidate, which missed the first name in the foreign term and is incorrect according to experts.
3.1.2 Translation lexical similarity
Translation lexical similarity uses automated translation of the foreign term and measures a lexical distance between that and the candidate terms. The distance of choice is the Levenshtein Levenshtein (1966) distance which computes how many changes are needed for the terms to be identical. We use a ratio with respect to length to normalize the distance across candidates.
Example: suppose the foreign term is “London School of Economics and Political Science". Automatic translation returns
$t=$
Gloss: "London School of Economics and Political Science."
We have the candidate:
$s=$
Gloss: "London School of Economics and Political Science."
The Levenshtein distance between $s$ and $t$ is 15. The ratio is calculated as
$\frac{\text{len}(t)+\text{len}(s)-15}{\text{len}(t)+\text{len}(s)}=0.79$
3.1.3 Transliteration lexical similarity
This feature captures the cases where the target Arabic term is just a transliteration of the source term by computing the Levenshtein ratio between the transliteration of the source term and the candidate. This is helpful in the case of translating proper nouns that should supposedly map phonetically.
Example: suppose the foreign term is "Ehud Prawer". After transliterating it, we get
$t=$
Gloss: "Ehud Brower"
We also have the candidate with a hamza instead of an alef
$s=$
Gloss: "Ehud Brower"
The Levenshtein distance between $s$ and $t$ is 1. The ratio is calculated as
$\frac{\text{len}(t)+\text{len}(s)-1}{\text{len}(t)+\text{len}(s)}=0.95$
3.1.4 Entity
A named entity is a real-world object or concept that can be clearly identified and categorized with a proper name.
Candidates that are detected as named entities have a higher probability of being a term. We use an open source Arabic NER model  Jarrar et al. (2022) to detect entities in Arabic context. Also, NER is used on the source term to favor candidates having the same entity type; we used the spaCy  Honnibal et al. (2020) model xx_ent_wiki_sm in particular.
3.1.5 Part of speech
A part of speech (POS) is a category of words with similar grammatical properties. Common parts of speech include nouns, verbs, adjectives, adverbs, pronouns, prepositions, conjunctions, and interjections. Terms are mostly noun phrases and rarely start with verbs.
We used POS tagging tools Obeid et al. (2020) with its maximum likelihood expectation morphological disambiguator to detect the POS of the first word occurring in a candidate.
3.1.6 Phonetic similarity
As illustrated in Table 3, we use the soundex algorithm Russell and Odell (1918) to determine whether the source and candidate terms sound the same. This serves a similar purpose to transliteration lexical similarity; however, practicaly characterized with higher precision and lower recall.
Table 3: Phonetic similarity examples
Source term Candidate Phonetically similar
Regavim Gloss: "Regavim" True
Lockheed Martin Gloss: "Lockheed Regavim" True
Lockheed Martin Gloss: "like Lockheed Regavim" False
4 Related work
Besides bilingual dictionaries, efforts have been deployed to develop lexical sources from several original languages into Arabic. Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields. Other more specialized glossaries have been developed such as the work of Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy . Furthermore, glossaries offering support to interpreters Sharif have also been established.
The Arabic digital ontology Jarrar (2022) includes digitized Arabic dictionaries as denoted by dictionary authors. It requires API access for term lookup.
However, these are not complete, need to be updated manually regularly for novel terms, and are not easy to integrate in the editing process.
5 Results
Table 4 shows statistics of the Arabic books we have processed and the annotated dataset we have built to train and evaluate our work.
Table 4: Statistics from processed books
Arabic books 496
Unique foreign terms extracted 58,300
All occurrences of foreign terms 83,516
Total candidates 301,326
Average number of candidates per occurrence 3.61
Number of annotated books 9
Number of foreign terms in annotated dataset 3,595
Number of candidates in annotated dataset 15,865
5.1 Heuristics
In this approach, we came up with the following formula after multiple iterations of observing the results and fine-tuning:
$s=S_{L}+S_{S}+S_{E}+S_{P}+S_{POS}$
1. Lexical score $S_{L}$:
$S_{L}=\mathbb{1}\left(l_{1}\geq\tau\right)\cdot(1.2\cdot l_{1})+\mathbb{1}\left(l_{2}\geq\tau\right)\cdot l_{2}$
where $l_{1}$ is the lexical similarity with the translation, $l_{2}$ is with the transliteration, and $\tau=0.7$ is the similarity threshold.
2. Semantic score $S_{S}$:
$S_{S}=1.45\cdot S$
where $S$ is the semantic similarity score.
3. Entity score $S_{E}$:
$S_{E}=\begin{cases}0.7,&e\neq\text{"none"}\text{ and }e=e_{s}\\ 0.3,&e\neq\text{"none"}\text{ and }e\neq e_{s}\\ 0,&\text{otherwise}\end{cases}$
where $e$ is the candidate’s entity label and $e_{s}$ is the source term’s entity label.
4. Phonetic score $S_{P}$:
$S_{P}=\begin{cases}1,&\text{if phonetically similar}\\ 0,&\text{otherwise}\end{cases}$
5. Part of speech score $S_{POS}$:
$S_{POS}=\begin{cases}-1,&\text{if pos}\in\{\text{"verb"},\text{"prep"},\text{"conj"}\}\\ 1,&\text{if pos}\in\{\text{"noun"},\text{"noun\_prop"}\}\\ 0,&\text{otherwise}\end{cases}$
$\mathbb{1}(A)$ is the indicator function:
$\mathbb{1}(A)=\begin{cases}1,&\text{if }A\text{ is true}\\ 0,&\text{otherwise}\end{cases}$
After that, we select the candidates (of the same occurrence of a foreign term) with the highest scores within a maximum deviation (taken 0.03 here). Table 5 shows the results achieved by the heuristics approach on the whole dataset. Note that fine-tuning was done only by observing the training data. These decent results provided a good starting point that later helped in the labeling process of the data.
5.2 Machine learning
An instance of the dataset is the features of a candidate, and a label that is True if the candidate is the target candidate and False otherwise.
We also augmented the feature space with the following: semantic similarity rank within the candidates related to the same occurrence of a foreign term, semantic similarity difference with the candidate ranked $1^{st}$, translation lexical similarity rank, translation lexical similarity difference, transliteration lexical similarity rank, transliteration lexical similarity difference. The intuition behind the augmentation is to provide comparative context to the ML model with respect to other candidates.
Let $x\in$semantic similarity, translation lexical similarity, transliteration lexical similarity},
$A_{i}^{x}$ = Array of values of feature $x$ of candidates related to the same occurrence $i$ of a term, sorted by decreasing order.
For each candidate j:
[…]
where the candidate j belongs to the occurrence occ(j).
As shown in Table 4, We annotated 15,865 candidates to build a training dataset. Then we trained a binary classifier using the Auto-WEKA package on the Weka Workbench Frank et al. (2016). We allowed Weka to run for several hours on 8 threads, which resulted in the following:
{strip}
Best classifier: weka.classifiers.trees.RandomForest
Arguments: [-I, 131, -K, 28, -depth, 0]
Attribute search: weka.attributeSelection.GreedyStepwise
Attribute search arguments: [-C, -R]
Attribute evaluation: weka.attributeSelection.CfsSubsetEval
Attribute evaluation arguments: [-M]
Metric: errorRate
Table 5 shows the results.
We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction instead of fixing a threshold, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False. This achieved a 92% precision and 98.4% recall on the total dataset, which is worse than the 99% precision and 98.5% recall achieved by the direct application of the Random Forest on the total dataset.
Table 5: Comparison of Heuristics and ML Results
Category Metric Training Testing
Heuristics Precision 87%
Recall 89%
F1 Score 88%
ML Precision 100% 94.9%
Recall 100% 92.4%
F1 Score 100% 93.7%
Dataset Size 12,692 (80%) 3,173 (20%)
5.3 Discussion and error analysis
Upon further inspection, it was observed that most of the classification errors made by the ML model involved cases that required expert judgment to discern their correctness. This highlights an advantage of the model: its errors occur primarily on inherently difficult cases, which are challenging even for experts. Table 6 shows an example.
Table 6: Selecting a wrong candidate for the foreign term SA (Gloss: "Sturmabteilung")
Candidate Prediction Label
Gloss: "Storm Battalion" False True
Another cause of errors are peculiar cases where there is one letter difference between the correct translation and the selected candidate, which is caused by Arabic prepositions linked to a word. This is usually handled through pre-processing or providing both candidates to the model for it to decide on one, which might fail in some cases. Table 7 shows an example.
Table 7: Selecting a wrong candidate for the foreign term Verteilungsstellen (Gloss: "distribution points")
Candidate Prediction Label
Gloss: "distribution entities" True True
Gloss: "by distribution entities" True False
6 Conclusion
This paper introduced methods for extracting terminology mappings between Arabic and foreign languages from Arabic books. The approach achieves useful results, contributing to automated Terminology Management and text processing. The created dataset and existing corpora further enrich the research community, advancing cross-lingual knowledge exchange.
Limitations
Some terms have subtle translations that are inherently complex and challenging to identify, even for human experts. Therefore, the tool is designed to assist rather than replace expert judgment, effectively handling simpler terms automatically while supporting experts with more difficult cases.
Ethics statement
The data was collected and used with the appropriate approvals of the intellectual property owners. All results are reported following best academic standards and practices.

### v2

@@
-Towards Terminology Management Automation for Arabic
+MASRAD: Arabic Terminology Management Corpora with Semi-Automatic Construction
 Abstract
-This paper presents a method and supporting tools for automation of terminology management for Arabic.
-The tools extract lists of parallel terminology matching terms in foreign languages to their Arabic counterparts from field specific texts.
-This has significant implications as it can be used to improve consistent translation and use of terms in specialized Arabic academic books, and provides automated aid for enhancing cross-lingual text processing.
-This automation of terminology management aims to reduce processing time, and ensure use of consistent and correct terminology.
-The extraction takes advantage of naturally occurring term translations. It considers several candidate phrases of varying lengths that co-occur next to the foreign terms.
-Then it computes several similarity metrics, including lexicographic, phonetic, morphological, and semantic ones to decide the problem.
-We experiment with heuristic, machine learning, and ML with post processing approaches.
-This paper reports on a novel curated dataset for the task, an existing expert reviewed industry parallel corpora, and on the performance of the three approaches.
-The best approach achieved $94.9\%$ precision and $92.4\%$ recall.
-Towards Terminology Management Automation for Arabic
+This paper presents MASRAD, a terminology dataset for Arabic terminology management, and a method with supporting tools for its semi-automatic construction.
+The entries in MASRAD are $(f,a)$ pairs of foreign (non-Arabic) terms $f$, appearing in specialized, academic and field-specific books next to their Arabic $a$ counterparts.
+MASRAD-Ex systematically extracts these pairs as a first step to construct MASRAD.
+MASRAD helps improving term consistency in academic translations and specialized Arabic documents, and automating cross-lingual text processing.
+MASRAD-Ex leverages translated terms organically occurring in Arabic books, and considers several candidate pairs for each term phrase.
+The candidate Arabic terms occur next to the foreign terms, and vary in length.
+MASRAD-Ex computes lexicographic, phonetic, morphological, and semantic similarity metrics for each candidate pair, and uses heuristic, machine learning, and machine learning with post-processing approaches to decide on the best candidate.
+This paper presents MASRAD after thorough expert review and makes it available to the interested research community.
+The best performing MASRAD-Ex approach achieved 90.5% precision and 92.4% recall.
+MASRAD: Arabic Terminology Management Corpora with Semi-Automatic Construction
 Mahdi Nasser  and Laura Sayyah  and Fadi A. Zaraket Arab Center for Research and Policy Studies, Doha {mnasser,lsayyah,fzaraket} @dohainstitute.edu.qa
 1 Introduction
-Translation of academic literature is essential to knowledge transfer.
-Linguistic and structural complexities exist.
-These affect consistency and clarity of translated material and may hinder knowledge accessibility.
-Terminology management concerns the complexities involved in (i) specifying the term that should be used upon the introduction of a new concept to a language, and (ii) and enforcing the use, hence (iii) ensuring consistency and correctness.
-The first and most important step is to build TermBase, a large database of parallel terms.
-Authors can utilize TermBase to enhance their writing processes when faced with a new concept; be it a foreign term to be translated to Arabic.
-Recent advances in natural language processing and computational linguistics have paved the way for automated systems that assist in cross-lingual text processing.
+Terminology management concerns (i) specifying terms used upon the introduction of novel concepts to a language, and (ii) enforcing the use, hence (iii) ensuring consistency and correctness.
+Translation from foreign languages to Arabic sources significant terminology inconsistency due to the diversity of translation experts and the lack of reference parallel term corpora.
+Recent systems like TURJUMAN Nagoudi et al.
+(2022) have improved Arabic machine translation quality, but do not focus on terminology alignment or consistency.
+Linguistic and structural complexities affect consistency and clarity of translations and may hinder access to knowledge.
+The first and most important step is to build a termbase and provide it as a reference to Arabic scholars and experts.
+We present MASRAD, a large dataset of parallel terms with pairs of foreign terms and their corresponding Arabic counterparts, sourced into use in professional, academic and scholastic Arabic contexts.
+We make MASRAD available for the research community and to authors looking for consistent Arabic terms when introducing a novel concept to Arabic.
+Recent advances in natural language processing (NLP) and computational linguistics (CL) have paved the way for automated systems that assist in cross-lingual text processing.
 However, existing systems, including LLMs, often struggle with the nuances of terminology in cross-lingual specific expert fields, where precision is paramount.
-This paper introduces a novel methodology and a supporting framework to extract terminology mappings from existing Arabic books, a rich source of terms when building a TermBase.
-When a foreign term occurs in Arabic books, usually between parentheses, it is preceded by its Arabic translation.
+This paper constructs MASRAD, and presents a novel methodology and a supporting framework (MASRAD-Ex) to extract and curate parallel foreign-Arabic terms from existing Arabic books.
+When a foreign term occurs in Arabic scholarly writing, it usually occurs between parentheses, and it is usually preceded by its Arabic translation.
 Two problems arise.
-The exact scope of the Arabic term is ambiguous, and often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones.
-Editors and reviewers, who are usually expert researchers, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
-Our work helps automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
-The problem is formulated as follows: given a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$, denote candidate terms as the set of strings $s_{i}=w_{1}\ldots w_{i}$ ending before term $f$, $1\leq i\leq n$ as a candidate term.
-The goal is to identify the target candidate term $s_{t}$ best matching $f$.
-Our tool extracts linguistic features for each candidate match $(s_{i},f)$.
-The tool interpolates the features and aggregates their values into a resulting score that ranks the candidate terms.
-By addressing the challenges of terminology mapping, this research moves us closer to a more integrated community where language barriers are significantly reduced, facilitating a greater exchange of knowledge and ideas across languages and cultures.
-We make verified parts of the constructed dataset available online for the research community.
-To the best of our knowledge, this approach to TermBase extraction for Arabic is novel, and no comparable datasets currently exist.
+(i) The exact scope of the Arabic term is ambiguous, and (ii) often times authors (even the same author) use Arabic terms inconsistently for the same non-Arabic ones.
+Editors and reviewers, who are usually expert researchers themselves, are required to manually solve these laborious and time consuming tasks during lengthy editing iterations.
+MASRAD and MASRAD-Ex help automate these two tasks to reduce review and edit efforts and help reviewers and editors concentrate on more important research and knowledge generation tasks.
+The problem is formulated as follows.
+Consider a sequence $s=\langle(f)w_{1}w_{2}w_{3}\ldots w_{n}\rangle$ representing a sentence with a foreign term $f$ between parentheses, and preceded from right to left with a sequence of Arabic words $w_{1},\ldots w_{n}$.
+We denote the candidate terms possible to match $f$ as the set of strings $S(f)=\{s_{i}:s_{i}=w_{1}\ldots w_{i},i\in[1,n]\}$ ending before term $f$.
+MASRAD-Ex extracts these candidates from the books, and identifies the target candidate term $s_{t}\in S(f)$ best matching $f$.
+MASRAD-Ex extracts linguistic features for each candidate match $(s_{i},f)$.
+It interpolates the features and aggregates their values into a resulting score, and then ranks the candidate terms.
+This work addresses the challenge of terminology mapping and reduces language barriers, facilitating a greater exchange of knowledge and ideas across languages and cultures.
+We make MASRAD and the constructed dataset of feature vectors (used for training/testing) available online for the research community upon request [^1] https://github.com/mnasser-dru/MASRAD.
+To the best of our knowledge, this approach to termbase building for Arabic is novel, and no comparable datasets currently exist.
 1.1 Motivating example
-Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme with different translations.
+Consider the following two contexts. They contain the same underlined french term l’ethnocentrisme (The ethnocentrism) with different translations.
   • Context 1:
@@
     Gloss: "He strongly criticized the ethnocentric tendency (l’ethnocentrisme) that seeks to replace primitive human ideas about the world and society with our European ideas."
-Candidates for Context 1:
+MASRAD-Ex computers the following candidates per context. Candidates for Context 1:
   • - Gloss: "And his sharp criticism of ethnocentrism does not turn into…"
@@
   • We propose a novel method for terminology mapping that leverages naturally occurring term translations, combining NLP and data driven techniques.
-  • We create a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
-  • We evaluate and compare three approaches: heuristic, machine learning, and machine learning with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges.
-The results are impressive and convincing for deployment.
-The final model runs locally, requires no LLM interaction, and thus requires taking no IP risks from Arabic publishing houses.
-We discuss TermBase creation in Section 2, the methodology in Section 3, briefly cover related work in  4, present and discuss our results in Section 5. We then discuss future directions and conclude.
+  • We build and provide MARSAD, a unique annotated dataset for terminology extraction, which can serve as a valuable resource for the research community.
+  • We build and evaluate MASRAD-Ex including three approaches: heuristic, machine learning (ML), and ML with post-processing demonstrating the effectiveness of these methods in addressing terminology mapping challenges.
+The results are impressive and industry/deployment level, convincing for deployment. The final model runs locally, requires no LLM interaction, and thus requires no data exposure or IP risks.
+We discuss MASRAD in Section 2, the methodology in Section 3, present and discuss results in Section 4, and cover related work in 5, We then discuss future directions and conclude.
 2 Dataset creation
-The source dataset is derived specifically from books published by our publishing house.
+Table 1: Statistics of all processed books
+Arabic books 495
+Unique foreign terms 58,570
+All occurrences of foreign terms 84,242
+Total candidates 334,564
+Average number of candidates per occurrence 3.97
+Table 2: Statistics of MASRAD (annotated candidates)
+Arabic books 15
+Unique foreign terms 4,347
+All occurrences of foreign terms 4,841
+Total candidates 19,405
+The source dataset is derived specifically from books published by a set of publishing houses that we keep anonymous for blind review.
 These books span multiple academic fields, offering a diverse range of terminologies.
 The dataset includes terms in foreign languages (usually enclosed in parentheses) alongside their surrounding Arabic text.
+Table 1 summarizes the data extracted from all books.
+Table 2 summarizes the data in MASRAD, our labeled and expert reviewed data that will be used later for training MASRAD-Ex.
 The labeling process involved extracting candidate Arabic terms preceding each foreign term and annotating them with a binary label: True if the candidate is the correct translation of the foreign term, and False otherwise.
-To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed to correct the errors.
+To streamline this process, a draft of labels was generated using heuristic-based rules, and were subsequently reviewed by experts to correct.
 2.1 Source data selection
@@
 Each extracted candidate term, along with its associated foreign term, was presented as an individual instance. The features of this instance, extracted using NLP tools, are described in section 3.
-Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 5.1 to facilitate the labeling process. This draft was then reviewed to ensure the accuracy of the labels.
+Subsequently, a draft of the labels was generated using the heuristics detailed in subsection 4.1 to facilitate the labeling process. This draft was then reviewed to correct errors.
 3 Methodology
@@
 3.1.1 Semantic similarity
-We utilized the sentence transformer LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
+We utilized the sentence transformer Reimers and Gurevych (2019) LaBSE Feng et al. (2022) to compute the embeddings of the source term and candidate term.
 After that, we compute semantic similarity as the cosine similarity between these 2 vectors Salton et al.
@@
 We used POS tagging tools Obeid et al. (2020) with its maximum likelihood expectation morphological disambiguator to detect the POS of the first word occurring in a candidate.
+An alternative is to use the POS tagging tools by Al‑Shaibani (2021).
 3.1.6 Phonetic similarity
@@
 Lockheed Martin Gloss: "like Lockheed Regavim" False
-4 Related work
-Besides bilingual dictionaries, efforts have been deployed to develop lexical sources from several original languages into Arabic.
-Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields.
-Other more specialized glossaries have been developed such as the work of Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy .
-Furthermore, glossaries offering support to interpreters Sharif have also been established.
-The Arabic digital ontology Jarrar (2022) includes digitized Arabic dictionaries as denoted by dictionary authors. It requires API access for term lookup.
-However, these are not complete, need to be updated manually regularly for novel terms, and are not easy to integrate in the editing process.
 4 Results
-Table 4 shows statistics of the Arabic books we have processed and the annotated dataset we have built to train and evaluate our work.
-Table 4: Statistics from processed books
-Arabic books 496
-Unique foreign terms extracted 58,300
-All occurrences of foreign terms 83,516
-Total candidates 301,326
-Average number of candidates per occurrence 3.61
-Number of annotated books 9
-Number of foreign terms in annotated dataset 3,595
-Number of candidates in annotated dataset 15,865
 4.1 Heuristics
@@
 $\mathbb{1}(A)=\begin{cases}1,&\text{if }A\text{ is true}\\ 0,&\text{otherwise}\end{cases}$
-After that, we select the candidates (of the same occurrence of a foreign term) with the highest scores within a maximum deviation (taken 0.03 here).
+After that, we select the candidate with the maximum score within an occurrence of a foreign term.
 Table 7 shows the results achieved by the heuristics approach on the whole dataset.
@@
 For each candidate $j$:
-[…]
-[…]
+$\text{x\_rank}(j)=A_{\text{occ}(j)}^{x}.\text{index\_of}(x(j))$
+$\text{x\_difference}(j)=\max(A_{\text{occ}(j)}^{x})-x(j)$
 where the candidate j belongs to the occurrence occ(j).
-As shown in Table 4, We annotated 15,865 candidates to build a training dataset.
+As shown in Table 1, We annotated 19,405 candidates to build the dataset.
 Then we trained a binary classifier using the Auto-WEKA package on the Weka Workbench Frank et al.
 (2016).
-We allowed Weka to run for several hours on 8 threads, which resulted in the following:
-{strip}
-Best classifier: weka.classifiers.trees.RandomForest
-Arguments: [-I, 131, -K, 28, -depth, 0]
-Attribute search: weka.attributeSelection.GreedyStepwise
-Attribute search arguments: [-C, -R]
-Attribute evaluation: weka.attributeSelection.CfsSubsetEval
-Attribute evaluation arguments: [-M]
+We allowed Weka to run for several hours on 8 threads, which resulted in the arguments displayed in Table 6.
+Table 6: AutoML arguments
+Best classifier: RandomForest
+Arguments: [-I, 252, -K, 0, -depth, 0]
+Attribute search: GreedyStepwise
+Attribute search arguments: [-C, -B, -R]
+Attribute evaluation: CfsSubsetEval
+Attribute evaluation arguments: [-L]
 Metric: errorRate
-Table 5 shows the results.
-We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction instead of fixing a threshold, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False.
-This achieved a 92% precision and 98.4% recall on the total dataset, which is worse than the 99% precision and 98.5% recall achieved by the direct application of the Random Forest on the total dataset.
+Table 7 summarizes the results.
+The training set comprises data from nine books, and the test set comprises data from six books.
+We split the data at the book level to make the task more challenging, where the model is tested on books it has never seen before, which better reflects a real-world scenario.
+We also tried ranking candidates by partitioning the testing dataset into groups of candidates related to the same occurrence of a term, and ranking them based on the value of their prediction, and then the candidate with the highest rank was chosen as the target term, and the others where detected as False.
+This achieved a 90.1% precision and 92.9% recall.
 Table 7: Comparison of Heuristics and ML Results
 Category Metric Training Testing
-Heuristics Precision 87%
-Recall 89%
-F1 Score 88%
-ML Precision 100% 94.9%
+Heuristics Precision 88.8% 81.5%
+Recall 88.2% 84.4%
+F1 Score 88.5% 82.9%
+ML Precision 100% 90.5%
 Recall 100% 92.4%
-F1 Score 100% 93.7%
-Dataset Size 12,692 (80%) 3,173 (20%)
+F1 Score 100% 91.4%
+Dataset Size 15865 3540
 4.3 Discussion and error analysis
@@
 Gloss: "by distribution entities" True False
+4.4 Evaluation against a manual glossary
+Over 495 books, our extraction pipeline identified 58,570 unique foreign terms.
+Intersecting these with our manual in‐house glossary of 46,787 concepts (a concept here consists of an English, a French, and an Arabic term) yielded 7,368 matches (15.74% of the glossary).
+This low overlap highlights a major problem in terminology management: assuming in the worst case that a concept was matched with both the English and French foreign terms of our extracted data, we get that at least 43,834 (74.84%) of the extracted terms are novel and need to be handled.
+Focusing on the 7,368 matched concepts, we evaluated our ranking model’s ability to surface the expert Arabic term.
+The correct glossary entry appeared as the top automated suggestion 30.6% of the time (top-1 accuracy), within the top two automated suggestions 32.99% of the time (top-2 accuracy), and within the top three automated suggestions 33.3% of the time (top-3 accuracy).
+By observing these accuracies, we can assume that the percentage of manual Arabic terms in these matched concepts that are also present in the 495 processed books to be close to our top-3 accuracy (33.3%).
+This reasoning is due to the minimal jumps between the top-k accuracies.
+From this we deduce two key insights:
+  • About 66.7% of the matched concepts have a different Arabic translation by the authors or translators of the processed books than that of the expert.
+This greatly highlights the inconsistency problem in terminology.
+  • Among the matched concepts with a present Arabic term in the books, our model placed it first in 91.89% of the cases (top-1 accuracy ÷ top-3 accuracy).
+This indicates that our model effectively elevates the expert translation when it is available.
+Note that the model performs really well when applying it to problems since it aggregates and compares results from the candidates of all occurrences of a term.
+5 Related work
+TURJUMAN Nagoudi et al.
+(2022) is a neural machine translation toolkit that translates into Modern Standard Arabic using fine-tuned AraT5 models.
+While effective at sentence-level translation, it does not address terminology alignment or consistency in specialized texts.
+In addition to bilingual dictionaries, efforts have been made to develop lexical sources from several original languages into Arabic.
+Arabic WordNet Global WordNet Association for instance relies on several open and restricted domain sources to develop a lexical English-Arabic database in different fields.
+Other more specialized glossaries have been developed, such as Hossam Mahdi’s Glossary of Terms for the Conservation of Cultural Heritage in Arabic Alphabetical Order Mahdy .
+Furthermore, glossaries offering support to interpreters Sharif have also been established.
+The Arabic digital ontology Jarrar (2022) includes digitized Arabic dictionaries as denoted by dictionary authors. It requires API access for term lookup.
+However, these are not complete, need to be updated manually regularly for novel terms, and are not easy to integrate in the editing process.
 6 Conclusion
-This paper introduced methods for extracting terminology mappings between Arabic and foreign languages from Arabic books.
-The approach achieves useful results, contributing to automated Terminology Management and text processing.
-The created dataset and existing corpora further enrich the research community, advancing cross-lingual knowledge exchange.
-Limitations
+This paper introduced the semi-automatic construction of MASRAD using MASRAD-Ex, a set of tools for extracting and aligning terminology.
+Our work achieves useful results, contributing to automated terminology management and text processing.
+The created dataset and corpora further enrich the research community, advancing cross-lingual knowledge exchange.
+Limitations and future work
 Some terms have subtle translations that are inherently complex and challenging to identify, even for human experts.
 Therefore, the tool is designed to assist rather than replace expert judgment, effectively handling simpler terms automatically while supporting experts with more difficult cases.
+As shown in Section 4.4, 84.26% of the entries in the manual glossary were not matched.
+This highlights the richness of the world of terminology and the necessity of drawing on additional sources to expand our termbase.
+To this end, we plan to apply MASRAD-Ex to newly acquired books and expand our tooling to support other type of sources.
+To improve the semantic similarity measure 3.1.1, we can fine-tune a sentence transformer Reimers and Gurevych (2019), or train a small model that augments the space of an existing one.
 Ethics statement